OpenStudy (anonymous):

Using Logarithms how would I go about solving the following? 1+ 2^x = 8 I currently have: log1 + 2logx = log8

OpenStudy (anonymous):

I can get 2logx = log8 - log1 but it doesn't make sense to me, please help. Thanks.

OpenStudy (wolf1728):

2^x = 7 then, taking logs of both sides x * log 2 = log 7

OpenStudy (jack1):

dunno about the logs dude, but logic would indicate that 1 + 7 = 8 so 2^x = 7 so x = log 7/ log 2

OpenStudy (jack1):

ah, my bad, wolf has faster internets ;D

OpenStudy (anonymous):

you guys are both right! its just x = log7/log2. Thanks guys! I wish I can give out two medals :(

OpenStudy (wolf1728):

Thanks zeno

OpenStudy (anonymous):

Thank you Wolf, I'll find something more challenging next time ;-)

OpenStudy (wolf1728):

That was challenging enough